QUESTION 20 In an M/M/1 when the arrival rate is the same as the service rate the number of customers waiting in line is zeroDelivery flexibility is a competitive dimension of operations strategy: True False
Answer #1

1.

Correct Answer:

False

As per the formula of the number of customers waiting in line is infinite or undefined.

For example, if service rate = arrival rate = 5 customers per hour

Then,

number of customers waiting in line = 5/(5-5) = 5/0 or undefined

============

2.

Correct answer:

False

It is the delivery reliability and process flexibility among the others, that work as competitive dimensions of the operations strategy.